### v1

@@ 4.5 Ablation Study
MovieLens 20M @1
0.049
0.060
0.054
0.059
@5
0.142
0.172
0.164
0.173
@10
0.211
0.251
0.243
0.252
Google Review @1 0.021 0.021
0.015
0.021
@5
0.060
0.061
0.052
0.061
@10
0.082
0.086
0.071
0.082
Amazon Review @1
0.038
0.034
0.037
0.041
@5
0.047
0.042
0.047
0.051
@10
0.050
0.047
0.051
0.055
@@
Table 7: Experiments on short-term text prompt long-term embedding. TP10: TextPrompt baseline with $10$ recent user events. TP50: TextPrompt baseline with $50$ recent user events. AR50: With $50$ user events as encoder inputs. AR50TP10: With $50$ user events as encoder inputs and $10$ most recent events in text format as LLM inputs.
Long-term and short-term user context The design of User-LLM allows it to integrate both long-term and short-term user context. We can feed long-term user history into the user encoder to generate compact user embeddings while prompting LLM with short-term user interactions. As shown in Table 7, User-LLM generally achieves better results when combining long-term and short-term user context for next item prediction tasks. Using 50 user activities for embedding generation and 10 most recent activities for prompting LLM, User-LLM (AR50TP10) outperforms TextPrompt with only short-term context (TP10) by providing long-term user information and outperforms TextPrompt with long-term context (TP50) by extracting meaningful user representations efficiently. Compared to AR50, in which User-LLM only uses long-term user data, short-term context provides the model (AR50TP10) with additional natural language information about item names. This leads to improved next item name prediction while preserving computational efficiency.
Benefit of pretraining We study the benefits of pretraining user encoders for downstream tasks. Results presented in columns 3 and 4 in Table 8 demonstrate that models utilizing pretrained encoders consistently outperform those with randomly initialized encoders across all evaluated tasks. This finding suggests that pretraining enables the encoder to capture latent user preference effectively, providing valuable context to the LLM during the cotraining process.
Soft Prompt v.s. Cross Attention We evaluate the performance of an alternative encoder-LLM integration strategy, soft-prompt Lester et al. (2021), within User-LLM framework, in which we prepend user embeddings as soft prompt tokens to the language model and meanwhile also learn an additional task prompt. We set the task prompt’s length to $10$ and utilized a pretrained Autoregressive encoder in our experiment. As shown in the last column of Table 8, the cross-attention-based approach generally outperforms the soft-prompt approach. This advantage is particularly significant in the two review generation tasks, suggesting that the cross-attention mechanism enables LLMs to better extract and utilize the rich user information encoded within user embeddings. This highlights the value of cross-attention for tasks demanding a nuanced understanding of human intent.
Dataset Task User-LLM
cross-attention soft prompt DualEnc -Xatten
PT RD
MovieLens 20M NextItem 0.054 0.041 0.051 0.044
FavGenre 0.785 0.778 0.787 0.775
Google Review NextItem 0.015 0.015 0.013 0.014
FavCat 0.862 0.853 0.822 0.797
ReviewG 11.72 11.43 9.06 11.68
Amazon Review NextItem 0.037 0.026 0.031 0.020
FavCat 0.890 0.882 0.894 0.854
ReviewG 26.38 23.12 25.80 23.83
Table 8: Experiments on different model fusing architectures, encoders, and the benefit of pretraining. Report Recall@1 for next item prediction and favorite genre/category prediction, ROUGE-Lsum for review generation. Full finetuning is used throughout the experiments. PT: Pretrained encoder. RD: Randomly initialized encoder. soft-prompt: Connect user embeddings with LLM via soft-prompt. DualEnc-Xatten: Use a dual encoder as the user encoder and cross-attend it to LLM.
Different Encoders We experiment with an alternative Transformer-based encoder, Dual Encoder (which serves as a baseline in this work), within the User-LLM framework. This architecture features two separate user and label towers to generate embeddings. The similarity between user and label embeddings is used for loss computation via a softmax layer during pretraining. We only the use user tower when integrating with LLMs. User-LLM with Dual Encoder performed worse than with AR encoder (Table 8). One reason is that Dual Encoder only outputs a single embedding for each user, providing less context to LLM compared to the AR encoder. Additionally, while Autoregressive Transformer benefits from loss calculation across all input tokens, Dual Encoder relies on a single label per input sequence. This limitation becomes more pronounced on sparse datasets like Amazon Review, leading to a larger performance gap.
5 Conclusion and Future Work
In this paper, we introduced User-LLM, a framework for contextualizing LLMs through user embeddings. These embeddings, derived from self-supervised pretraining on diverse user interactions, capture hidden user preferences and their evolution. By integrating these embeddings with LLMs through cross-attention and soft-prompt, User-LLM empowers LLMs to adjust dynamically to user contexts.
Our comprehensive evaluation across MovieLens, Amazon Review, and Google Local Review datasets demonstrated significant performance improvements in various tasks. User-LLM showed competitive performance compared with non-LLM baselines and text-prompt-based LLM personalization techniques, particularly in handling long sequences and understanding users deeply. User-LLM’s computational efficiency and ability to preserve LLM knowledge further make it a highly suitable approach for real-world user understanding applications.
Future research could delve into optimizing user embedding generation through advanced pretraining techniques, investigating the alignment between user embeddings and the language model space for deeper user context understanding, and training User-LLM on a diverse range of tasks to enhance its generalization abilities and adaptability across a broader spectrum of user scenarios. By addressing these aspects, User-LLM has the potential to become an even more robust and versatile framework for user modeling and LLM personalization.
Appendix A Encoder Architectures
User-LLM uses a Transformer-based user encoder to generate user embeddings from ID-based feature sequences. To effectively fuse information from multimodal user interaction data, we explore and compare two fusion architectures: early fusion and late fusion. Following the example shown in Fig. 4, each item in the user interaction timeline has three modalities: name, rating, and category. Each modality (or feature) has its own vocabulary. Items from each modality are mapped to an integer ID and have their own embedding representations. In an early fusion architecture, the $i_{th}$ item’s three modalities (name, rating, and category) are combined into a fused embedding $f_{i}$. The sequence of fused embeddings is then fed into the user encoder to generate user embeddings. On the other hand, a late fusion architecture employs three separate feature encoders to process each modality independently. The resulting embeddings are combined to a single user embedding at a later stage using fusion layers.
@@ Appendix A Encoder Architectures
Autoregressive Transformer As described in Section 3.1, User-LLM employs an Autoregressive Transformer as the early fusion encoder. Illustrated in Fig. 2, this model receives a sequence of user activities, where each activity is represented by multiple modalities (name, rating, and category). Individual features are first embedded, then concatenated and processed by a Transformer decoder. The Autoregressive Transformer Encoder generates a user embedding for each input item in the user’s timeline.
Compared to Autoregressive Transformers, Dual Encoders offer several advantages. The separate feature encoders within user towers allow independent processing of features, handling features with varying sequence lengths or misaligned timestamps and potentially enhancing information extraction. The configurable fusion layers provide flexibility in the number of generated embeddings. Additionally, Dual Encoders support diverse pretraining tasks beyond next item prediction, as user and label towers can process different features (e.g., name for the user tower and category for the label tower). However, Dual Encoders may be less effective at modeling the complex sequential patterns often present in multimodal user data. The optimal encoder choice depends on dataset characteristics, the desired level of contextual understanding, and a balance between computational cost and potential performance gain. Refer to Table 8, Table 12, and Table 13 for performance comparison between the two encoder architectures.
Appendix B Integration Approaches
We explored the integration of user embeddings and LLMs in User-LLM through two approaches: cross-attention and soft-prompt. In the cross-attention method, the output embeddings from the pretrained user encoder are cross-attended with the LLM’s intermediate text representations, similar to how Flamingo Alayrac et al. (2022) works. On the other hand, the soft-prompt approach leverages user embeddings from the pretrained encoder as a soft prompt for the LLM. This soft prompt is prepended to the regular text prompt embedding, providing additional context about the user’s interests and preferences. For a comparison of cross-attention and soft-prompt approaches, refer to Table 8 and Table 14.
Appendix C Training Hyperparameters
Dataset Task Learning rate Weight decay Batch size Training steps
MovieLens 20M
EncoderPretrain
1e-1
1e-1
32,768
10,000
NextItem
1e-3
1e-3
8,192
10,000
FavGenre
1e-3
1e-3
1,024
5,000
Google Review
EncoderPretrain
1e-4
1e-1
16,384
10,000
NextItem
1e-3
1e-1
2,048
5,000
FavCat
1e-3
1e-1
2,048
5,000
ReviewG
1e-3
1e-1
1,024
5,000
Amazon Review
EncoderPretrain
1e-4
1e-1
12,800
7,000
NextItem
1e-3
1e-1
8,192
10,000
FavCat
1e-3
1e-1
1,024
5,000
ReviewG
1e-3
1e-1
2,048
5,000
Table 9: Training hyperparameters.
Appendix D More Results
D.1 Next Item Prediction
Table 11 and Table 10 show results comparing text-prompt-based approach and User-LLM in next item prediction tasks.
Dataset Recall TP User-LLM
Google Review @1 0.021 0.015
@5 0.061 0.050
@10 0.086 0.071
Amazon Review @1 0.034 0.037
@5 0.042 0.047
@10 0.047 0.051
Table 10: User-LLM v.s. TextPrompt(TP) baseline for next item prediction on Google Local Review and Amazon Review datasets. Note that we use full finetune for both User-LLM and TP.
Recall Unfrozen LLM Frozen LLM
Len50 Len100 Len200 Len50 Len100 Len200
@1 0.060 0.054
0.034
0.037
0.018
0.020
0.023
0.053
0.017
0.040
0.011
0.031
@5 0.172 0.164
0.101
0.118
0.058
0.067
0.071
0.164
0.053
0.127
0.034
0.102
@10 0.251 0.243
0.151
0.181
0.090
0.106
0.108
0.244
0.080
0.198
0.053
0.159
D.2 Autoregressive Encoder v.s. Dual Encoder
@@ D.1 Next Item Prediction
MovieLens 20M @1 0.054 0.059
0.044
0.051
@5 0.164 0.173
0.135
0.151
@10 0.243 0.252
0.206
0.221
Google Review @1
0.015
0.021
0.014
0.020
@5
0.052
0.061
0.046
0.055
@10
0.071
0.082
0.066
0.078
Amazon Review @1 0.037 0.041
0.020
0.030
@5 0.047 0.051
0.024
0.038
@10 0.051 0.055
0.026
0.040
@@
D.3 Training Strategies and Encoder Pretraining
Dataset Metric AREnc-Xatten(User-LLM) DualEnc-Xatten
Full LoRA Enc Proj Full + PT
PT RD PT RD PT RD PT RD
NextItem MovieLens20M
R@1
0.054
0.041
0.053
0.043
0.052
0.041
0.040
0.002
0.044
R@5
0.164
0.133
0.164
0.136
0.164
0.133
0.134
0.008
0.142
R@10
0.243
0.203
0.244
0.205
0.243
0.204
0.203
0.013
0.218
Google Review
R@1
0.015 0.015 0.016 0.015 0.015 0.015 0.015 0.015 0.014
R@5
0.050
0.052 0.051
0.050
0.051 0.051
0.050
0.042
0.046
R@10
0.071 0.071
0.070
0.070
0.070
0.070
0.070
0.057
0.066
Amazon Review
R@1
0.037
0.023
0.028
0.018
0.028
0.022
0.014
0.000
0.020
R@5
0.047
0.031
0.035
0.026
0.034
0.029
0.019
0.000
0.024
R@10
0.051
0.037
0.038
0.029
0.036
0.030
0.021
0.0003
0.026
FavGenre (FavCat) MovieLens20M
R@1
0.785
0.778
0.784
0.779
0.783
0.778
0.750
0.484
0.775
R@5
0.996
0.994
0.996
0.994
0.996
0.994
0.993
0.931
0.994
R@10
0.999
0.999
0.999
0.999
0.999
0.999
0.999
0.991
0.999
Google Review
R@1
0.862 0.853
0.844
0.785
0.844
0.775
0.615
0.142
0.797
R@5
0.951 0.944
0.938
0.915
0.937
0.911
0.855
0.505
0.920
R@10
0.964 0.956
0.948
0.941
0.949
0.936
0.909
0.660
0.944
Amazon Review
R@1
0.890
0.882
0.887
0.884
0.885
0.886
0.850
0.223
0.854
R@5
0.934 0.948
0.917
0.931
0.917
0.930
0.918
0.663
0.912
R@10
0.944 0.958
0.919
0.934
0.919
0.932
0.923
0.793
0.934
ReviewG
Google Review
ROUGE
11.72
11.43
11.64
11.55
11.76
11.56
9.61
8.05
11.68
Amazon Review
ROUGE
26.38
23.12
24.56
25.00
24.30
25.18
19.04
17.30
23.83
D.4 Cross-attention v.s. Soft-prompt
@@ D.3 Training Strategies and Encoder Pretraining
FavGenre (FavCat) MovieLens20M R@1 0.785 0.784 0.783 0.750 0.787 0.787 0.781 0.692
@@
Appendix E Review Generation Examples
Table 15 shows some examples from Amazon/Google review generation tasks. From these examples, we can observe personalized responses generated using User-LLM exhibit greater alignment with user preferences and backgrounds compared to unpersonalized responses. While discrepancies remain between personalized responses using User-LLM and the ground truth, User-LLM demonstrates the ability to produce reviews that reflect specific user characteristics and inclinations more effectively.
Input Prompt Ground Truth Response Personalized Response Unpersonalized Response
“Write a review for The Firm - Aerobic Body Shaping.”
“Save Your Money.”
“Terrible Waste of Time.”
“Aerobic Workout.”
“Write a review for A.I. - Artificial Intelligence.”
“A pure movie masterpiece of a film by Steven Speilberg. You will either love or hate this movie, but it will make you think.”
“If you are into science fiction and you are looking for a movie that is not only good but also educational and fun, ’A.I.”
“A.I. - Artificial Intelligence.”
“Write a review for The Family that Preys.”
“Four Stars.”
“Four Stars.”
“A great thriller!”
“Write a review for Green Street Hooligans 2.”
“Hmmmmmm…”
“Not for everyone.”
“A good sequel.”
“Help me write a review for Blue with a rating of 5.0.”
“(Translated by Google) Very delicious and the service staff are very enthusiastic (Original) 非常好吃,服务人员很热忱.”
“(Translated by Google) Good food (Original) 好吃.”
“Great food and service.”
“Help me write a review for The Home Depot with a rating of 2.0.”
“(Translated by Google) The service is terrible (Original) El servicio esta pésimo.”
“(Translated by Google) Very bad service (Original) Muy mal servicio.”
“No help.”
Table 15: Review generation examples.

### v2

@@ Appendix D Integration Approaches
MovieLens 20M @1 0.049 0.049 0.055 0.059
@5 0.142 0.140 0.154 0.173
@10 0.211 0.206 0.243 0.252
Google Review @1 0.021 0.021 0.015 0.021
@5 0.060 0.061 0.052 0.061
@10 0.082 0.086 0.071 0.082
Amazon Review @1 0.038 0.034 0.037 0.041
@5 0.047 0.042 0.047 0.051
@10 0.050 0.047 0.051 0.055
@@
Table 8: Experiments on short-term text prompt plus long-term embedding. TP10: TextPrompt baseline with $10$ recent user events. TP50: TextPrompt baseline with $50$ recent user events. AR50: With $50$ user events as encoder inputs. AR50TP10: With $50$ user events as encoder inputs and $10$ most recent events in text format as LLM inputs.
@@
Dataset Metric AREnc-Xatten(User-LLM) DualEnc-Xatten
Full LoRA Enc Proj Full Enc Proj
PT RD PT RD PT RD PT RD PT RD PT RD PT RD
Next Item MovieLens 20M R@1 0.055 0.041 0.051 0.043 0.050 0.041 0.023 0.002 0.044 0.025 0.034 0.010 0.032 0.001
R@5 0.154 0.133 0.164 0.136 0.164 0.133 0.134 0.008 0.142 0.102 0.116 0.043 0.107 0.008
R@10 0.243 0.203 0.244 0.205 0.243 0.204 0.203 0.013 0.218 0.169 0.182 0.076 0.166 0.012
Google Review R@1 0.015 0.015 0.016 0.015 0.015 0.015 0.015 0.015 0.014 0.015 0.014 0.015 0.015 0.015
R@5 0.052 0.052 0.051 0.050 0.051 0.051 0.050 0.042 0.046 0.044 0.047 0.049 0.049 0.044
R@10 0.071 0.071 0.070 0.070 0.070 0.070 0.070 0.057 0.066 0.063 0.061 0.067 0.064 0.057
Amazon Review R@1 0.037 0.023 0.028 0.018 0.028 0.022 0.014 0.000 0.020 0.005 0.016 0.002 0.008 0.000
R@5 0.047 0.031 0.035 0.026 0.034 0.029 0.019 0.000 0.024 0.010 0.019 0.005 0.010 0.002
R@10 0.051 0.037 0.038 0.029 0.036 0.030 0.021 0.0003 0.026 0.014 0.019 0.009 0.011 0.003
Fav Genre/Category MovieLens 20M R@1 0.787 0.778 0.787 0.779 0.786 0.778 0.743 0.484 0.775 0.704 0.776 0.759 0.649 0.484
R@5 0.996 0.994 0.996 0.994 0.996 0.994 0.993 0.931 0.994 0.959 0.993 0.980 0.976 0.931
R@10 0.999 0.999 0.999 0.999 0.999 0.999 0.999 0.991 0.999 0.987 0.999 0.997 0.998 0.991
Google Review R@1 0.862 0.853 0.844 0.785 0.844 0.775 0.615 0.142 0.797 0.642 0.760 0.508 0.271 0.142
R@5 0.951 0.944 0.938 0.915 0.937 0.911 0.855 0.505 0.920 0.842 0.885 0.711 0.607 0.505
R@10 0.964 0.956 0.948 0.941 0.949 0.936 0.909 0.660 0.944 0.871 0.914 0.798 0.763 0.661
Amazon Review R@1 0.890 0.882 0.887 0.884 0.885 0.886 0.850 0.223 0.854 0.745 0.843 0.714 0.312 0.223
R@5 0.934 0.948 0.917 0.931 0.917 0.930 0.918 0.663 0.912 0.870 0.898 0.866 0.694 0.662
R@10 0.944 0.958 0.919 0.934 0.919 0.932 0.923 0.793 0.934 0.925 0.930 0.903 0.846 0.793
Review Gen Google Review Rouge 11.72 11.43 11.64 11.55 11.76 11.56 9.61 8.05 11.71 11.04 11.76 10.82 9.18 6.62
Amazon Review Rouge 26.38 23.12 24.56 25.00 24.30 25.18 19.04 17.30 23.83 18.1 21.11 20.78 18.92 18.53
@@
MovieLens 20M @1 0.055 0.059 0.044 0.051
@5 0.164 0.173 0.135 0.151
@10 0.243 0.252 0.206 0.221
Google Review @1 0.015 0.021 0.014 0.020
@5 0.052 0.061 0.046 0.055
@10 0.071 0.082 0.066 0.078
Amazon Review @1 0.037 0.041 0.020 0.030
@5 0.047 0.051 0.024 0.038
@10 0.051 0.055 0.026 0.040
@@
FavGenre (FavCat) MovieLens20M R@1 0.787 0.787 0.786 0.743 0.787 0.787 0.781 0.692
@@
Appendix E More Results
Recall Unfrozen LLM Frozen LLM
Len50 Len100 Len200 Len50 Len100 Len200
@1 0.049 0.055 0.043 0.055 0.034 0.055 0.017 0.051 0.016 0.051 0.015 0.053
@5 0.140 0.154 0.123 0.153 0.102 0.155 0.051 0.147 0.046 0.146 0.042 0.146
@10 0.206 0.227 0.183 0.226 0.154 0.229 0.077 0.214 0.072 0.214 0.066 0.216
E.1 Autoregressive Encoder v.s. Dual Encoder
E.2 Training Strategies and Encoder Pretraining
E.3 Cross-attention v.s. Soft-prompt
E.4 Next Item Prediction
Table 12 show results comparing text-prompt-based approach and User-LLM in next item prediction task on MovieLens20M dataset with varying input activity sequence length.
E.5 Long-term and short-term user context
User-LLM can integrate both long-term and short-term user context. We can feed long-term user history into the user encoder to generate compact user embeddings while prompting LLM with short-term user interactions. As shown in Table 8, User-LLM generally achieves better results when combining long-term and short-term user context for next item prediction tasks. Using 50 user activities for embedding generation and 10 most recent activities for prompting LLM, User-LLM (AR50TP10) outperforms TextPrompt with only short-term context (TP10) by providing long-term user information and outperforms TextPrompt with long-term context (TP50) by extracting meaningful user representations efficiently. Compared to AR50, in which User-LLM only uses long-term user data, short-term context provides the model (AR50TP10) with additional natural language information about item names. This leads to improved next item name prediction while preserving computational efficiency.
Appendix F Review Generation Examples
We presented some examples from Amazon/Google review generation tasks below. From these examples, we can observe personalized responses generated using User-LLM exhibit greater alignment with user preferences and backgrounds compared to unpersonalized responses. While discrepancies remain between personalized responses using User-LLM and the ground truth, User-LLM demonstrates the ability to produce reviews that reflect specific user characteristics and inclinations more effectively.
  • Input prompt: “Write a review for The Firm - Aerobic Body Shaping.”
    Ground truth response: “Save Your Money.”
    Personalized response: “Terrible Waste of Time.”
    Unpersonalized response: “Aerobic Workout.”
  • Input prompt: “Write a review for A.I. - Artificial Intelligence.”
    Ground truth response: “A pure movie masterpiece of a film by Steven Speilberg. You will either love or hate this movie, but it will make you think.”
    Personalized response: “If you are into science fiction and you are looking for a movie that is not only good but also educational and fun, ’A.I.”
    Unpersonalized response: “A.I. - Artificial Intelligence.”
  • Input prompt: “Write a review for The Family that Preys.”
    Ground truth response: “Four Stars.”
    Personalized response: “Four Stars.”
    Unpersonalized response: “A great thriller!”
  • Input prompt: “Write a review for Green Street Hooligans 2.”
    Ground truth response: “Hmmmmmm…”
    Personalized response: “Not for everyone.”
    Unpersonalized response: “A good sequel.”
  • Input prompt: “Help me write a review for Blue with a rating of 5.0.”
    Ground truth response: “(Translated by Google) Very delicious and the service staff are very enthusiastic.”
    Personalized response: “(Translated by Google) Good food.”
    Unpersonalized response: “Great food and service.”
  • Input prompt: “Help me write a review for The Home Depot with a rating of 2.0.”
    Ground truth response: “(Translated by Google) The service is terrible (Original) El servicio esta pésimo.”
    Personalized response: “(Translated by Google) Very bad service (Original) Muy mal servicio.”
    Unpersonalized response: “No help.”